OpenStudy (anonymous):

express in simplest form: 1 - 1/x --------- x - 2 + 1/x

OpenStudy (anonymous):

\[\frac{1-\frac{1}{x}}{x-2+\frac{1}{x}}\] like that?

OpenStudy (anonymous):

yes like that

OpenStudy (anonymous):

clear the compound fraction by multiplying numerator and denominator by \(x\), takes one step

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

multiply top and bottom by \(x\)

OpenStudy (anonymous):

\[\frac{1-\frac{1}{x}}{x-2+\frac{1}{x}}\times \frac{x}{x}\] is the first step

OpenStudy (anonymous):

you are allowed to do this because \(\frac{x}{x}=1\) it is just a trick to clear the denominators

OpenStudy (anonymous):

so its gonna be x−1 --------- x^2−2x+1

OpenStudy (anonymous):

the you factor ? 9x-1)(x-1)

OpenStudy (anonymous):

(x-1) *

OpenStudy (anonymous):

\[\frac{x-1}{x^2-2x+1}=\frac{x-1}{(x-1)(x-1)}=\frac{1}{x-1}\] i guess that is right